# Adversarial Style Augmentation via Large Language Modelfor Robust Fake News Detection

### v2

@@ A.1 Data & Implementation Details
Table 4: Performance comparison of different attack scenarios on the PolitiFact Dataset. (A: Adversarial prompt, B: Summarization prompt, C: In-Context prompt D: Adversarial Paraphraser)
@@
Comparison on other possible attack scenarios.
@@
We have conducted additional comparison experiments with more diverse attack scenarios: (1) Adversarial prompt: Given a news article and its label, the prompt instructs the LLM to rewrite the article to evade detection as the given label. (2) Summarization prompt: A prompt instructing the LLM to summarize the news article without incorporating stylistic elements. (3) In-Context prompt: A prompt providing an example of a recent, real CNN article, instructing the LLM to rewrite the given article in the same style. (4) Adversarial Paraphraser: The attack utilizes a paraphraser adversarially trained on the training dataset. The formats for each attack prompt are described in Figures 9 to 11. Based on the experimental results in Table 4, we confirm that our proposed model demonstrates a non-trivial performance improvement compared to other baselines against these attacks.
@@
<EXAMPLES>: [news article] Determine whether the given <EXAMPLES> content is fact or fake, mentioning the style of the article that led to that decision following format: Answer: {Fact or Fake} Reason: {one-sentence reason }
@@
Your current article is: [news article] But this article is detected as [Answer] due to the following [Reason]. Based on the above information, rewrite a new improved article not to be detected as [news label], maintaining the original content, as follows:
@@
Figure 9: The format of Adversarial prompt. First, use the prompt above to extract the answer and reason of factuality of the given news article. Then, use the following prompt for paraphrasing to avoid detection based on the identified reason. The "news article" section contains the original news text, and the "news label" section contains the corresponding news label.
@@
Figure 10: The format for Summarization prompt. The “news article" part will be filled with the original news text.
<EXAMPLES>: [news article example] Rewrite the following article as the writing style of <EXAMPLES> : [news article]
Figure 11: The format for In-Context prompt. The "news article example" section is filled with an example from a specific publisher, while the "news article" section contains the original news text.
Comparison with LLM-based Baselines.
We compared AdStyle with other LLM-based baselines: GPT-3.5 OpenAI (2022), GPT-4 OpenAI (2023), and In-Context Learning Brown et al. (2020). OpenAI APIs are utilized to predict news article labels, with the temperature set to 0 to ensure stable and consistent predictions. In the case of In-context Learning, one example each of fake news and real news was provided. The prompt formats for the LLM-based baselines are as follows:
Figure 12: Prompt for LLM-based baselines. The "news article" section contains the original news text.
Table 5: Performance comparison of different LLM-based baselines on the PolitiFact Dataset. (NY: The New York Times, TS: The Sun, NE: National Enquirer)
Model CNN NY TS NE Clean
GPT-3.5 0.5820 0.6242 0.6173 0.5274 0.7037
GPT-4 0.7133 0.6748 0.7025 0.6021 0.7823
In-Context 0.6954 0.6504 0.6875 0.5754 0.7383
Ours 0.9174 0.9444 0.9520 0.9460 0.9460
Full result on style-conversion attack.
Table 6 reports the AUC for each style-conversion attack. Paired sample t-tests were conducted to compare our method with the second-best baseline, Sheepdog, across different ratios for each dataset. Except for the 25% ratio in the Gossipcop dataset, all p-values were significantly smaller than 0.05, indicating statistically significant differences.
Table 6: Performance comparison with AdStyle in four different style—conversion attacks. For the attack scenario, we report the average AUC of four style-conversion attacks. The best results are marked in bold. Our model demonstrates a significant performance improvement over all text-based fake news detectors in diverse style-conversion attack scenarios. (NY: The New York Times, TS: The Sun, NE: National Enquirer )
CNN Politifact Gossipcop Constraint
0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
Vanilla 0.5664 0.6566 0.8127 0.6841 0.7913 0.8518 0.8563 0.8175 0.8056
UDA 0.5906 0.7174 0.8009 0.7294 0.7859 0.8599 0.8305 0.8594 0.9021
RADAR 0.7205 0.7301 0.7926 0.7683 0.8138 0.8666 0.8099 0.8400 0.8873
ENDEF 0.5426 0.6679 0.7588 0.7288 0.7839 0.8624 0.7695 0.8666 0.8353
SheepDog 0.6305 0.8269 0.8683 0.7472 0.8393 0.8706 0.9200 0.9022 0.9544
AdStyle 0.7547 0.8822 0.9174 0.8195 0.8429 0.8744 0.9344 0.9430 0.9646
NY Politifact Gossipcop Constraint
0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
Vanilla 0.5899 0.7334 0.8687 0.6750 0.7797 0.8465 0.8522 0.8939 0.8868
UDA 0.6189 0.7644 0.8744 0.7207 0.7795 0.8558 0.8745 0.9184 0.9297
RADAR 0.6966 0.7853 0.8586 0.7481 0.7960 0.8574 0.8701 0.8974 0.9176
ENDEF 0.6152 0.7630 0.8150 0.7227 0.7837 0.8609 0.8341 0.9067 0.8918
SheepDog 0.6310 0.7990 0.9090 0.7420 0.8309 0.8650 0.8937 0.9254 0.9645
AdStyle 0.7599 0.8940 0.9444 0.8118 0.8389 0.8686 0.9243 0.9594 0.9720
TS Politifact Gossipcop Constraint
0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
Vanilla 0.6913 0.6859 0.8789 0.7068 0.7918 0.8430 0.7822 0.8727 0.8974
UDA 0.6646 0.8072 0.8883 0.7488 0.7904 0.8613 0.8662 0.9145 0.9446
RADAR 0.7271 0.7173 0.8909 0.7594 0.8062 0.8620 0.8564 0.8882 0.9132
ENDEF 0.7001 0.8132 0.8444 0.7603 0.7909 0.8639 0.8445 0.8955 0.9043
SheepDog 0.7070 0.8408 0.9096 0.7580 0.8395 0.8688 0.8962 0.9217 0.9665
AdStyle 0.8444 0.9002 0.9520 0.8117 0.8347 0.8751 0.9219 0.9521 0.9727
NE Politifact Gossipcop Constraint
0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
Vanilla 0.6432 0.6856 0.8591 0.7020 0.7875 0.8399 0.7833 0.8853 0.9067
UDA 0.6557 0.7892 0.8622 0.7450 0.7902 0.8595 0.8747 0.9191 0.9423
RADAR 0.7428 0.7269 0.8864 0.7573 0.8026 0.8556 0.8775 0.8919 0.9161
ENDEF 0.6924 0.7877 0.8355 0.7502 0.7897 0.8589 0.8456 0.9112 0.9025
SheepDog 0.6636 0.8268 0.9166 0.7521 0.8331 0.8632 0.8606 0.9257 0.9667
AdStyle 0.8029 0.8913 0.9460 0.8166 0.8391 0.8702 0.9092 0.9579 0.9771

### v3

@@ 4.3. Performance Analysis
Table 6. Comparison under attack scenarios on the PolitiFact Dataset (A: Adversarial prompt, B: Summarization prompt, C: In-Context prompt D: Adversarial Paraphraser).
@@
(a) Diversity
(b) Adversarialness
(c) Coherency
Figure 7. Qualitative analyses with y-axis representing (a) diversity in embedding ($\mathbf{z}_{c}$) of conversion prompts sampled by AdStyle and adversarial-only selection; (b) Adversarialness ($s_{adv}^{c}$) for selected and unselected prompts; and (c) Coherency for selected and unselected prompts.
Figure 8. Comparison of augmented samples generated by our model and SheepDog: (A) Histogram of lexical diversity, (B) Histogram of perplexity, and (C) Perplexity across rounds.
Effectiveness of the selection strategy. We here empirically verified that our selection strategy (Sec. 3.3) effectively selects diverse style-conversion prompts with high adversarialness and coherency. Figure 7(a) visualizes the diversity of prompts selected by our method compared to those chosen solely based on adversarialness scores (i.e., Adversarial-only Selection, the third model in our ablation study). We measure diversity using the average cosine similarity of every pair of selected prompts’ embeddings, $\mathbf{z}_{c}$ (Eq. 2):
(Eq. 4) $\text{Diveristy}(\mathcal{C})$ $=1-{1\over|\mathcal{C}|}\sum_{(c_{i},c_{j})\in\mathcal{C}}{\text{sim}(\mathbf{z}_{c_{i}},\mathbf{z}_{c_{j}})},$
Round 0
P1: Rewrite the following article in a nonsensical and absurdly exaggerated tone with a hint of horror
P2: Rewrite the following article in a sarcastic and mocking tone
P3: Rewrite the following article in a chaotic and disorganized tone
Round 1
P1: Rewrite the following article in a haunting and macabre tone with a sense of impending horror and madness
P2: Rewrite the following article in a cryptic and enigmatic tone
P3: Rewrite the following article in a malevolent and apocalyptic tone with a sense of impending doom and destruction, while also incorporating elements of surrealism and existential dread
Figure 9. Example adversarial style-conversion prompts selected for the PolitiFact dataset.
Figure 7(b) and 7(c) illustrate the Adversarialness and Coherency, respectively, of our selected prompts compared to remaining unselected prompts. Adversarialness was measured using the $s_{adv}^{c}$ score (Section 3.3), and coherency was calculated as the cosine similarity between the original text and its augmented version using semantic BERT embeddings (Chanchani and Huang, 2023). We can also observe that, for both metrics, prompts selected through AdStyle exhibit higher values compared to unselected prompts. This suggests that our strategy effectively selects prompts by considering both adversarialness and coherency.
Analysis on style-conversion prompts. Finally, we conducted a qualitative analysis of our style-conversion prompts to understand what characteristics of the augmented samples contribute to improving the detector’s robustness. Interestingly, when we compare the augmented samples from two models: AdStyle and SheepDog, we found that our model produced sentences with significantly higher perplexity according to the language model backbone used by the detector than SheepDog model (9.05 vs. 4.26, see Figure 8B), even though ours have lower lexical diversity (0.503 vs. 0.634, see Figure 8A). In other words, our generated samples featured sentence structures that the detector’s language model likely had not encountered during pretraining. By providing the detector with inputs that have a variety of sentence structures and styles it has not previously seen, the detector naturally becomes robust against style-conversion attacks. In addition, when we measured the changes in perplexity of the augmented samples across training rounds (see Figure 8C), the perplexity increased over rounds and eventually converged at a certain point. This indicates that iterative exploration over multiple rounds is effective in creating augmentations that increasingly challenge the detector.
Figure 9 shows examples of prompts generated and selected by AdStyle. In contrast, our prompts included creative phrases beyond typical human-crafted suggestions, making them adversarial to the detector. The round 0 result shows a prompt asking for “(P1) a nonsensical and absurdly exaggerated tone with a hint of horror”, while round 1 offers a lengthier prompt mentioning, for example, “(P3) a malevolent and apocalyptic tone with a sense of impending doom and destruction, while also incorporating elements of surrealism and existential dread.” Previous work relied on conventional prompts, such as “neutral” or “sensational.” The LLM is particularly well-suited for exploring such a vast array of candidates. Currently, we use a fixed format for the initial set of prompts, which results in the generation of prompts with a similar format. We expect that using a broader set of prompts will allow the LLM to optimize prompts within a wider search space.
5. Conclusion
We presented a robust fake news detection method that effectively withstands paraphrasing attacks through adversarial style conversion. Unlike traditional detectors that use predefined and agnostic augmentations, AdStyle employs tailored augmentations that shift samples in the direction of the detector’s current decision boundary using style-conversion prompts, functioning similarly to adversarial noise. Among the various candidates for the LLM prompts, we selected an efficient set of prompts for training by considering diversity, coherence, and adversarialness.
Consequently, we were able to train a detector that exhibits high robustness and generalizability against a wide range of attacks. This robustness ensures that the detector can effectively identify and mitigate various forms of style-conversion attacks, regardless of the target news outlet. We believe that our research, combined with the shared codes, contributes to better filtering of fake news online. We aim to support the broader goal of ensuring the integrity of information in the digital age.
Ethical Consideration
We recognize several ethical concerns in this work. First, our method may inherit biases from the training data, which can be mitigated by carefully selecting style-conversion results that preserve textual coherence. Second, while our approach aims to detect fake news, similar techniques could be misused to enhance misinformation—an ongoing challenge in adversarial research. Lastly, our detection methods might influence news production by constraining creative writing styles, highlighting the need for further research on the balance between algorithmic moderation and information systems.
Acknowledgments
J.G. Lee and M. Cha are co-corresponding authors. We extend our gratitude to Fangzhao Wu, Wenchao Dong, and the anonymous reviewers for their insightful feedback on our work. This research was supported by the National Research Foundation of Korea (NRF) grant (RS-2022-00165347).
Appendix A APPENDIX
A.1. Baseline Implementation
For consistency, experiments replicating existing baselines maintained fixed settings for learning, backbone network, and other relevant parameters. Augmented variations for UDA were generated using back-translation through German. RADAR employed adversarial training to learn paraphrases, utilizing the maximization of the binary cross-entropy loss of a fake news detector as a reward signal, with the T5-large model serving as the paraphraser. Named entities for ENDEF were extracted using the bert-base-NER model from Hugging Face. For SheepDog, augmentations were generated following the original paper’s methodology, using the prompt format illustrated in Figure 10 and four tones: ”objective and professional,” ”neutral,” ”emotionally triggering,” and ”sensational”.
Rewrite the following article in a/an [tone]: [news article]
Figure 10. Prompt for style-conversion. The “tone” part will be filled with the desired tone, and the “news article” part will contain the original news text.
A.2. Examples of Style-Conversion Prompt
Figure 11 showcases examples generated by the selected style-conversion prompt, demonstrating how fake news content can be modified while preserving its meaning. These examples demonstrate the impact of style transformation on fake news detection.
Original: Updated: ASPEN 2014 Senator John McCain is meeting with his top campaign advisers and donors here…
Rewrite the following article in a sarcastic and mocking tone: Oh, how exciting! Senator John McCain is gracing us with his presence in Aspen, meeting with his elite group of advisers and donors…
Rewrite the following article in a chaotic and disorganized tone: Senator John McCain is like, in Aspen, meeting with his campaign peeps and rich donors, you know?…
Rewrite the following article in a nonsensical and absurdly humorous tone with a touch of surrealism and a sprinkle of silliness : Senator John McCain found himself in a pickle. He was surrounded by his top campaign advisers and donors in the mystical town of Aspen…
Figure 11. Generated example input created using the selected style-conversion prompt.
A.3. Prompt Examples of Other Attack Scenarios
@@
We have conducted comparison experiments with more diverse attack scenarios: (1) Adversarial prompt: Given a news article and its label, the prompt instructs the LLM to rewrite the article to evade detection as the given label. (2) Summarization prompt: A prompt instructing the LLM to summarize the news article without incorporating stylistic guidance. (3) In-Context prompt: A prompt providing an example of a recent, real CNN article, instructing the LLM to rewrite the given article in the same style. Here we provide the example prompt for each attack prompt as figures. (4) Adversarial Paraphraser: The attack utilizes a paraphraser adversarially trained on the training dataset.
¡EXAMPLES¿: [news article]
@@
Determine whether the given ¡EXAMPLES¿ content is fact or fake, mentioning the style of the article that led to that decision following format:
Answer: {Fact or Fake}
Reason: {one-sentence reason}
Your current article is: [news article]
But this article is detected as [Answer] due to the following [Reason].
@@
Based on the above information, rewrite a new improved article not to be detected as [news label], maintaining the original content, as follows:
@@
Figure 12. Example input for Adversarial prompt model. First, use the prompt above to extract the answer and reason for the authenticity of the given news. Next, utilize the prompt below to rephrase the news to evade detection based on the given reason. The ”news article” section contains the original news text, and the ”news label” part contains the corresponding label.
@@
Figure 13. Example input for Summarization prompt model. The “news article” part is filled with the original news text.
¡EXAMPLES¿: [news article example] Rewrite the following article as the writing style of ¡EXAMPLES¿ : [news article]
Figure 14. Example input for In-Context prompt model. The ”news article example” section is filled with an example from a specific publisher, while the ”news article” section contains the original news text.